OpenStudy (anonymous):

√ 6 x + 5 = √ 53 PLEASE SHOW STEPS

OpenStudy (anonymous):

\[\sqrt{6x+5}=\sqrt{53}\] is this it?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

THEN WHAT

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

vijay has it right :)